### v1

# Dispersion of first sound in a weakly interacting ultracold Fermi liquid: an exact calculation

Abstract
At low temperature, a normal gas of unpaired spin-1/2 fermions is one of the cleanest realizations of a Fermi liquid. It is described by Landau’s theory, where no phenomenological parameters are needed as the quasiparticle interaction function can be computed perturbatively in powers of the scattering length $a$, the sole parameter of the short-range interparticle interactions. Obtaining an accurate solution of the transport equation nevertheless requires a careful treatment of the collision kernel, as the uncontrolled error made by the relaxation time approximations increases when the temperature $T$ drops below the Fermi temperature. Here, we study sound waves in the hydrodynamic regime up to second order in the Chapman-Enskog’s expansion. We find that the frequency $\omega_{q}$ of the sound wave is shifted above its linear depart as $\omega_{q}=c_{1}q(1+\alpha q^{2}\tau^{2})$ where $c_{1}$ and $q$ are the speed and wavenumber of the wave and the typical collision time $\tau$ scales as $1/a^{2}T^{2}$. Besides the shear viscosity, the coefficient $\alpha$ is described by a single second-order collision time which we compute exactly from an analytical solution of the transport equation, resulting in a positive dispersion $\alpha>0$. Our results suggest that ultracold atomic Fermi gases are an ideal experimental system for quantitative tests of second order hydrodynamics.
Introduction— Landau’s Fermi liquid theory is an effective theory which, when it is applicable, greatly simplifies the description of the many-body dynamics of a system of fermions, breaking it down to a single kinetic equation on a distribution of dilute quasiparticles. It is a very successful theory in describing the phenomenology of a wide class of fermionic systems, such as liquid 3He Baym and Pethick (1991); Vollhardt and Wolfle (1990), electron gases Jones and McClure (1966); Behnia (2022); Belitz and Kirkpatrick (2022), quantum gases Vichi (2000), down to nuclear/neutron matter Schäfer (2014). Nevertheless, once the phenomenological parameters characterizing the quasiparticle dispersion and interaction function are known, either from measurements Greywall (1986); Matsumoto et al. (1996) or from a microscopic calculation Lipschitz and Pitaevskii (1981a), Fermi liquid theory provides quantitative predictions on dynamical properties, such as the transport coefficients Højgård Jensen et al. (1968); Brooker and Sykes (1968).
In normal 3He, despite decades of research, there is still a discrepancy between theory and experiment on the value of the transport coefficients. Consider for instance the shear viscosity $\eta$, which scales with temperature as $T^{-2}$: theory still underestimate the product $\eta T^{2}$ from the measurement by more than $20\%$ Abel et al. (1966); Baym and Pethick (1991); Parpia, J. M. et al. (1978); González et al. (2016); Baten et al. (2023). This is due to a limited knowledge of the quasiparticle interaction function, which has not be been computed from a microscopic theory, and whose experimental determination is limited to the lowest spherical harmonics Baym and Pethick (1991). As a consequence, the exact solutions of the transport equation Højgård Jensen et al. (1968); Brooker and Sykes (1968) were never validated experimentally, and relaxation time approximations Khalatnikov and Abrikosov (110); Baym and Pethick (1991) remain in use today González et al. (2016).
Ultracold gases of fermionic gases provide exciting opportunities to quantitatively test those transport calculations Huang et al. (2024). These gases behave as Fermi liquids when the $s$-wave scattering length $a$ is negative and sufficiently small to open a regime of temperatures $T_{c}\ll T\ll T_{\textrm{F}}$ where $T_{c}$ is the superfluid critical temperature and $T_{\textrm{F}}$ the Fermi temperature. A Fermi liquid regime may also exist when $a$ is positive and small enough to suppress three-body recombination Ji et al. (2022). The quasiparticle dispersion and interaction function can be computed perturbatively in powers of $k_{\textrm{F}}|a|$ (with $k_{\textrm{F}}$ the Fermi wavenumber) Lipschitz and Pitaevskii (1981a). Experimentally, both the interaction strength and the temperature can be varied such that the typical collision time $\tau$ can be adjusted over several orders of magnitude Riedl et al. (2008); Patel et al. (2020), to explore both collisionless and hydrodynamic regimes Huang et al. (2024). Flat-bottom potentials Gaunt et al. (2013); Navon et al. (2021), where sound can be excited at very low wavevector $q$ in homogeneous samples allow to study the propagation and attenuation of sound waves in a very controlled environment Patel et al. (2020).
Theoretically, great efforts were devoted to the calculation of the viscosity at strong coupling, in particular in the unitary regime $|a|=+\infty$ Rupak and Schäfer (2007); Bruun and Smith (2007); Enss et al. (2011) and exact results are available in the high temperature virial regime Braby et al. (2010); Nishida (2019); Fujii and Enss (2023). At intermediate temperatures however, a controlled approach has not been found due to the absence of a separation of timescales between the collisional and kinetic dynamics Lipschitz and Pitaevskii (1981b). The temperature range of the Fermi liquid regime shrinks as the quasiparticle cross section increases with the interaction strength Pavaskar and Rothstein (2022), and it is eventually hidden by the onset of a superfluid phase at a critical temperature $T_{c}\approx 0.17T_{\textrm{F}}$ at unitarity. Below $T_{c}$, sound attenuation is dominated by phonon-phonon interactions Rupak and Schäfer (2007); Enss et al. (2011); Kurkjian et al. (2017). In the weakly interacting normal phase, the transport coefficients were computed using relaxation time approximations Vichi (2000); Massignan et al. (2005); Watabe et al. (2010); Maki and Enss (2023), (either in the “variational” Massignan et al. (2005); Maki and Enss (2023), or in the original Abrikosov-Khalatnikov Vichi (2000); Watabe et al. (2010) formulation), even though their uncontrolled error increases towards low temperatures. In this Letter, we perform an exact calculation of the transport coefficients Højgård Jensen et al. (1968); Brooker and Sykes (1968) to lowest order in $k_{\textrm{F}}|a|$, and show that the error of the relaxation time approximations is significant, up to 25%.
For negative values of $a$, the quasiparticle interactions are attractive Recati and Stringari (2010), which prevents the emergence of a zero sound mode as in liquid 3He. We thus lack a parameter similar to $c_{0}-c_{1}$, where $c_{0}$, $c_{1}$ are the speed of zero and first sound respectively, to characterize the dispersion of sound, as was done in 3He Abel et al. (1966). In this work, we derive the leading-order deviation of the frequency $\omega_{q}$ from its linear depart $c_{1}q$. To do so, we solve the transport equation to the order $\tau^{2}$ of the Chapman-Enskog’s expansion. This is the so-called second hydrodynamic order often used in a relativistic context to cure the acausality of the diffusion equations characteristic of dissipative hydrodynamics Israel (1976). We find the exact solution of the transport equation by decomposing the quasiparticle distribution function on a basis of orthogonal polynomials adapted to the low temperature limit Hofmann and Gran (2023). Remarkably, the frequency shift involves only two parameters of the collision kernel: the viscous relaxation time $\tau_{\eta}$ and a second-order viscous time $t_{\eta}$, which we both compute exactly. As for the dispersion of the sound branch Kurkjian et al. (2017), we find that $\omega_{q}$ is above its linear depart $c_{1}q$, the deviation being proportional to $q^{3}\tau^{2}$ with $\tau$ scaling as $1/a^{2}T^{2}$.
Transport equation— Landau’s theory postulates that a Fermi liquid is described by a local quasiparticle distribution $n_{\sigma}(\textbf{p},\textbf{r},t)$, which is the number of quasiparticles of spin $\sigma$ having momentum p at position r and time $t$. This distribution slightly deviates from its thermal equilibrium value $n^{\rm eq}_{\sigma}$ due to some slowly-varying (in both space and time) perturbation. The energy of an arbitrary quasiparticle configuration is expanded to second ordrer:
$E=E_{0}+\sum_{\textbf{p},\textbf{r},\sigma}\epsilon_{\rm eq,\sigma}(\textbf{p})\delta n_{\sigma}(\textbf{p},\textbf{r})\\ +\frac{1}{2}\sum_{\textbf{p},\sigma,\textbf{p}^{\prime},\sigma^{\prime},\textbf{r}}f_{\sigma\sigma^{\prime}}(\textbf{p},\textbf{p}^{\prime})\delta n_{\sigma}(\textbf{p},\textbf{r})\delta n_{\sigma^{\prime}}(\textbf{p}^{\prime},\textbf{r})$ (1)
where $\delta n_{\sigma}=n_{\sigma}-n^{\rm eq}_{\sigma}$ is the fluctuation about thermal equilibrium $n_{\sigma}^{\rm eq}(\epsilon_{\rm eq})=1/(1+\textrm{e}^{(\epsilon_{\rm eq}-\mu)/T})$. In the general case, the equilibrium dispersion relation $\epsilon_{\rm eq}$ and interaction function $f_{\sigma\sigma^{\prime}}$ are phenomenological parameters usually rexpressed in terms of an effective mass and Landau parameters. In the case of a weakly interacting gas with contact interactions, these quantities can be calculated perturbatively in powers of the coupling constant $g=4\pi a/m$ Lipschitz and Pitaevskii (1981a) (we use $\hbar=k_{\rm B}=1$ throughout this work). To first order in perturbation theory, we have:
$\epsilon_{\sigma}^{\rm eq}(\textbf{p})=\frac{p^{2}}{2m}+g\rho_{{\rm eq},\sigma^{\prime}},\quad f_{\uparrow\downarrow}={g}/V,\quad f_{\sigma\sigma}=0$ (2)
where $V$ is the volume of the gas.
The evolution of the quasiparticle distribution is described by a transport equation:
$\frac{\partial n_{\sigma}}{\partial t}+\frac{\partial\epsilon_{\sigma}}{\partial\textbf{p}}\cdot\frac{\partial n_{\sigma}}{\partial\textbf{r}}-\frac{\partial(\epsilon_{\sigma}+U_{\sigma})}{\partial\textbf{r}}\cdot\frac{\partial n_{\sigma}}{\partial\textbf{p}}=I_{\rm coll,\sigma}$ (3)
where $U_{\sigma}$ is an external driving field, $\epsilon_{\sigma}(\textbf{p},\textbf{r})=\epsilon_{\sigma}^{\rm eq}(\textbf{p})+\sum_{\textbf{p}^{\prime},\sigma^{\prime}}f_{\sigma\sigma^{\prime}}(\textbf{p},\textbf{p}^{\prime})\delta n_{\sigma^{\prime}}(\textbf{p}^{\prime},\textbf{r})$ is the local energy of the quasiparticles, and the collision integral is given in this weakly interacting limit by Fermi’s golden rule
$I_{\sigma}(\textbf{p})=\frac{2\pi g^{2}}{V^{2}}\sum_{\textbf{p}_{2},\textbf{p}_{3},\textbf{p}_{4}}\delta_{\textbf{p}+\textbf{p}_{2},\textbf{p}_{3}+\textbf{p}_{4}}\delta\left(\epsilon_{\sigma}(\textbf{p})+\epsilon_{\sigma^{\prime}}(\textbf{p}_{2})-\epsilon_{\sigma}(\textbf{p}_{3})-\epsilon_{\sigma^{\prime}}(\textbf{p}_{4})\right)\\ \left[(1-n_{\sigma}(\textbf{p}))(1-n_{\sigma^{\prime}}(\textbf{p}_{2}))n_{\sigma}(\textbf{p}_{3})n_{\sigma^{\prime}}(\textbf{p}_{4})-n_{\sigma}(\textbf{p})n_{\sigma^{\prime}}(\textbf{p}_{2})(1-n_{\sigma}(\textbf{p}_{3}))(1-n_{\sigma^{\prime}}(\textbf{p}_{4}))\right].$ (4)
In this expression, all the quasiparticle distributions $n$ are evaluated at position r and time $t$, which reflects the assumption that collisions are local and instantaneous.
As we seek the eigenmodes of the transport equation, we assume that the drive is weak and linearize Eq. (3) around equilibrium. We focus here on the unpolarized case so $\mu$ is the common chemical potential of the two spin species. We also restrict ourselves to excitations of the total density, and define the total driving field $U_{\rm tot}=U_{\uparrow}+U_{\downarrow}$, that we decompose in Fourier space $U_{\rm tot}(\textbf{r},t)=\text{Re}(U\sum_{\textbf{q}}\textrm{e}^{\textrm{i}(\textbf{q}\cdot\textbf{r}-\omega t)})$. In Fourier space, the transport equation obeyed by $\delta n=\delta n_{\uparrow}+\delta n_{\downarrow}$ becomes:
$\left(\omega-\frac{\textbf{p}\cdot\textbf{q}}{m}\right)\delta n(\textbf{p})+\frac{\partial n_{\rm eq}}{\partial\epsilon_{\rm eq}}\frac{\textbf{p}\cdot\textbf{q}}{m}\left(g\delta\rho+U\right)=\textrm{i}I_{\rm lin}$ (5)
where $\delta\rho(\textbf{q},\omega)=(1/V)\sum_{\textbf{p}^{\prime}}\delta n(\textbf{p}^{\prime},\textbf{q},\omega)$ is the fluctuation of the total density about $\rho_{\rm eq}=\rho_{\rm eq,\uparrow}+\rho_{\rm eq,\downarrow}$, and the linearized collision integral $I_{\rm lin}(\{\delta n(\textbf{p}^{\prime})\}_{\textbf{p}^{\prime}})=I_{\rm lin,\uparrow}+I_{\rm lin,\downarrow}$ is obtained by linearizing Eq. (4) with respect to the function $n_{\sigma}(\textbf{p})$.
For $T\ll T_{\rm F}$, transport occurs in a energy shell of typical depth $T$ around the Fermi energy $\epsilon_{\rm F}$ Nozières and Pines (1966); Baym and Pethick (1991). We thus reparametrize the quasiparticle distribution as
$\delta n(\textbf{p})=U\frac{\partial n_{\rm eq}}{\partial\epsilon_{\rm eq}}\nu(\epsilon,\theta)\text{ with }\epsilon=\frac{\epsilon_{\rm eq}-\mu}{T}=\frac{p^{2}/2m-\epsilon_{\textrm{F}}}{T}$ (6)
where we have used the equation of state $\mu=\epsilon_{\textrm{F}}+g\rho_{\rm eq}/2$ of the weakly interacting Fermi gas. Note that the reduced energy $\epsilon$ varies from $\infty$- ∞ to +$\infty$+ ∞ as the momentum $p$ varies from 00 to +$\infty$+ ∞.
Restricting the transport equation to momenta $p$ lying in the relevant energy shell around the Fermi surface, we obtain, to leading order in $T/T_{\textrm{F}}$:
$(c-\cos\theta)\nu(\epsilon,\theta)-\frac{k_{\textrm{F}}a}{\pi}\cos\theta\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}g(\epsilon^{\prime})\int_{0}^{\pi}\sin\theta^{\prime}\textrm{d}\theta^{\prime}\nu(\epsilon^{\prime},\theta^{\prime})\\ +\frac{\textrm{i}}{\pi\omega_{0}\tau}\left[\Gamma(\epsilon)\nu(\epsilon,\theta)+\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}\int\sin\theta^{\prime}\textrm{d}\theta^{\prime}\frac{\textrm{d}\phi^{\prime}}{2\pi}\mathcal{N}_{\rm od}(\epsilon,\epsilon^{\prime},u)\nu(\epsilon^{\prime},\theta^{\prime})\right]=-\cos\theta$ (7)
where $g(\epsilon)=1/(4\text{cosh}^{2}(\epsilon/2))$ is the dimensionless density of states, $\omega_{0}=v_{\textrm{F}}q$ is a typical excitation frequency (with $v_{\textrm{F}}=\sqrt{2\epsilon_{\textrm{F}}/m}$ the Fermi velocity), $c=\omega/\omega_{0}$ is the dimensionless excitation frequency, and
$\tau=\frac{1}{2ma^{2}T^{2}}$ (8)
is a typical collision time. The diagonal part of the collision kernel is given by
$\Gamma(\epsilon)=\pi^{2}+\epsilon^{2}.$ (9)
This dimensionless function sets the physical quasiparticle lifetime to $\tau_{\rm qp}(\epsilon)=\pi\tau/\Gamma(\epsilon)$. The off-diagonal part of the kernel reads
$\mathcal{N}_{\rm od}(\epsilon,\epsilon^{\prime},u)={\frac{\mathcal{S}(\epsilon,-\epsilon^{\prime})}{\sqrt{2(1+u)}}-2\frac{\mathcal{S}(\epsilon,\epsilon^{\prime})}{\sqrt{2(1-u)}}}\\ \text{with }\mathcal{S}(\epsilon,\epsilon^{\prime})=\frac{\epsilon-\epsilon^{\prime}}{2}\frac{\text{cosh}\frac{\epsilon}{2}}{\text{cosh}\frac{\epsilon^{\prime}}{2}\text{sinh}\frac{\epsilon-\epsilon^{\prime}}{2}}$ (10)
where the angular dependence comes through $u=\cos(\textbf{p},\textbf{p}^{\prime})=\cos\theta\cos\theta^{\prime}+\sin\theta\sin\theta^{\prime}\cos(\phi-\phi^{\prime})$.
The conservation of the number of quasiparticle in a collision provides a relation between $\Gamma$ and $\mathcal{S}$:
$\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}\mathcal{S}(\epsilon,\epsilon^{\prime})=\int_{-\infty}^{+\infty}\textrm{d}\epsilon^{\prime}\frac{g(\epsilon^{\prime})}{g(\epsilon)}\mathcal{S}(\epsilon^{\prime},\epsilon)=\frac{\Gamma(\epsilon)}{2}.$ (11)
The hydrodynamic limit— The hydrodynamic limit is the regime of short collision times
$\omega_{0}\tau\ll 1.$ (12)
In this regime, collisions bring the quasiparticle distribution back to equilibrium much faster than the typical time $1/\omega_{0}$ at which the sound wave evolves. Only the few components of the distribution that are not affected by collisions (i.e. those that belong to the zero-energy space of the collision kernel) remain significantly excited.
To obtain the exact solution of the transport equation in this limit, we expand the distribution $\nu$ over a basis of orthogonal polynomials
$\nu(\epsilon,\theta)=\sum_{n,l}\nu_{n}^{l}P_{l}(\cos\theta)Q_{n}(\epsilon)$ (13)
where $P_{l}$ are the Legendre polynomials, and $Q_{n}$ are orthogonal for the scalar product weigthed by the density of states:
$\int_{-\infty}^{+\infty}g(\epsilon)Q_{n}(\epsilon)Q_{m}(\epsilon)\textrm{d}\epsilon=||Q_{n}||^{2}\delta_{nm}.$ (14)
The polynomials $Q_{n}$ are obtain by the usual recurrence relation:
$\epsilon Q_{n}=Q_{n+1}+\xi_{n}Q_{n-1}\quad\text{with}\quad\xi_{n}\equiv\frac{||Q_{n}||^{2}}{||Q_{n-1}||^{2}}$ (15)
We choose $Q_{0}=1$ and $Q_{1}=\epsilon$ as the initial condition. Note that even and odd polynomials are respectively symmetric and antisymmetric about the Fermi surface: $Q_{n}(-\epsilon)=(-1)^{n}Q_{n}(\epsilon)$. By contrast to this exact approach, the relaxation time approximations Khalatnikov and Abrikosov (110); Baym and Pethick (1991) truncate the expansion in Eq. (13) to $n=0$, thereby neglecting the energy dependence of the quasiparticle distribution. The difference between the Abrikosov-Khalatnikov Khalatnikov and Abrikosov (110) and the variational formulation Baym and Pethick (1991); Massignan et al. (2005); Braby et al. (2010) lies in the treatment of the remaining $\epsilon$ dependence of the collision kernel: Abrikosov and Khalatnikov replaced it by its value in $\epsilon=0$ (in particular they approximate the quasiparticle lifetime by its value at the Fermi level $\Gamma(\epsilon)\approx\Gamma(0)$), while the variational formulation averages it over $\epsilon$.
Note that the polynomials $Q_{n}$ used here differ from the orthogonal polynomials of the momentum $p$ used at nonvanishing temperature Braby et al. (2010); Huang et al. (2024), and that the replacement Eq. (6) converts [^1] Denoting $H_{n}$ and $T_{n}$ the polynomials defined respectively by the first and second line of Eq. (S85) of the supplementary material of Ref. Huang et al. (2024) (so as to avoid the confusion with the $Q_{n}$ polynomials defined in this work), one has $Q_{n}(\epsilon)=\lim_{m\to+\infty}H_{2n}(\sqrt{\epsilon+m})=\lim_{m\to+\infty}\frac{H_{2n+1}(\sqrt{\epsilon+m})}{\sqrt{\epsilon+m}}=\lim_{m\to+\infty}\frac{T_{2n+2}(\sqrt{\epsilon+m})}{m}$ even/odd powers of $p^{2}$ into even/odd powers of $\epsilon$.
The matrix elements of the collision kernel in the orthogonal basis $\{Q_{n}\}$ are given by
$\Gamma_{nn^{\prime}}=\int_{-\infty}^{+\infty}\textrm{d}\epsilon g(\epsilon)\Gamma(\epsilon)\frac{Q_{n}(\epsilon)}{||Q_{n}||^{2}}Q_{n^{\prime}}(\epsilon)$ (16)
and
$\!\!\!\mathcal{N}_{nn^{\prime}}^{l}\!=\!\!\int_{-\infty}^{+\infty}\!\!\!\!\!\!\!\!\textrm{d}\epsilon^{\prime}\textrm{d}\epsilon g(\epsilon)\int_{-1}^{1}\!\!\!\!\textrm{d}uP_{l}(u)\frac{Q_{n}(\epsilon)}{||Q_{n}||^{2}}\mathcal{N}_{\rm od}(\epsilon,\epsilon^{\prime},u)Q_{n^{\prime}}(\epsilon^{\prime})\\ =\frac{2}{2l+1}\mathcal{S}_{nn^{\prime}}\left((-1)^{l+n^{\prime}}-2\right)$ (17)
with
$\mathcal{S}_{nn^{\prime}}=\int_{-\infty}^{+\infty}\textrm{d}\epsilon\textrm{d}\epsilon^{\prime}g(\epsilon)\mathcal{S}(\epsilon,\epsilon^{\prime})\frac{Q_{n}(\epsilon)}{||Q_{n}||^{2}}Q_{n^{\prime}}(\epsilon^{\prime}).$ (18)
Note that the subspaces of symmetric and antisymmetric functions of $\epsilon$ (even and odd $n$ respectively) are decoupled: $\Gamma_{nn^{\prime}}=\mathcal{N}_{nn^{\prime}}^{l}=0$ if ${n+n^{\prime}}$ is odd Brooker and Sykes (1970). This is specific to the low-temperature limit where both the density of state $g(\epsilon)$ and energy integration domain are symmetric about the Fermi surface. Treating separately the odd and even orders, $\Gamma$ and $\mathcal{S}$ are tridiagonal matrices and can be expressed analytically as
$\!\Gamma_{nn^{\prime}}\!\!\!$ $=$ $\!\!\!\left(\pi^{2}+\xi_{n+1}+\xi_{n}\right)\delta_{nn^{\prime}}+\delta_{n-2,n^{\prime}}+{\delta_{n+2,n^{\prime}}}\xi_{n+2}\xi_{n+1}$
$\mathcal{S}_{nn^{\prime}}\!\!\!$ $=$ $\!\!\!\left[2\pi^{2}\frac{n^{2}+n-1}{4n^{2}+4n-3}\right]\delta_{nn^{\prime}}+\frac{\delta_{n-2,n^{\prime}}}{n(n-1)}$ (19)
$\qquad\qquad\qquad\qquad+\frac{\delta_{n+2,n^{\prime}}\xi_{n+2}\xi_{n+1}}{(n+2)(n+1)}$
with $\xi_{n}=\frac{\pi^{2}n^{4}}{(2n+1)(2n-1)}$.
The transport equation projected on the orthogonal basis reads now
$c\nu_{n}^{l}-\left[\frac{l+1}{2l+3}\nu_{n}^{l+1}+\frac{l}{2l-1}\nu_{n}^{l-1}\right]-\frac{2k_{\textrm{F}}a}{\pi}\delta_{l,1}\delta_{n,0}\nu_{0}^{0}\\ +\frac{\textrm{i}}{\pi\omega_{0}\tau}\sum_{n^{\prime}}\mathcal{M}_{nn^{\prime}}^{l}\nu_{n^{\prime}}^{l}=-\delta_{l,1}\delta_{n,0}$ (20)
where we introduce the complete collision tensor
$\mathcal{M}_{nn^{\prime}}^{l}=\Gamma_{nn^{\prime}}+\mathcal{N}_{nn^{\prime}}^{l}.$ (21)
The conservation of the number of quasiparticles, energy, and momentum in a collision [see Eq. (4)] generates zero-energy eigenfunctions of the collision kernel. In our orthogonal basis, this simply translates into $\mathcal{M}_{0n^{\prime}}^{0}=\mathcal{M}_{n0}^{0}=0$, $\mathcal{M}_{1n^{\prime}}^{0}=\mathcal{M}_{n1}^{0}=0$ and $\mathcal{M}_{0n^{\prime}}^{1}=\mathcal{M}_{n0}^{1}=0$, respectively, The corresponding equations of motion on the conserved quantities $\nu_{0}^{0}$, $\nu_{0}^{1}$ and $\nu_{1}^{0}$ are
$c\nu_{0}^{0}-\frac{\nu_{0}^{1}}{3}$ $=$ $0,$ (22)
$c\nu_{0}^{1}-\left(1+\frac{2k_{\textrm{F}}a}{\pi}\right)\nu_{0}^{0}-\frac{2}{5}\nu_{0}^{2}$ $=$ $-1,$ (23)
$c\nu_{1}^{0}-\frac{\nu_{1}^{1}}{3}$ $=$ $0$ (24)
Physically $\nu_{0}^{0}=-{2\pi^{2}}\chi_{\rho}/mk_{\textrm{F}}$ is proportional to the density response $\chi_{\rho}=\delta\rho/U$, $\nu_{0}^{1}=-k_{\textrm{F}}\chi_{v}/2$ is proportional to the response of the velocity $\chi_{v}=v_{\parallel}/U$ with $\rho_{\rm eq}v_{\parallel}=({1}/{V})\sum_{\textbf{p}}(\textbf{p}\cdot\textbf{q}/mq)\delta n(\textbf{p})$, and $\nu_{1}^{0}=-{6}\chi_{e}/{mTk_{\textrm{F}}}$ is proportional to the response of the energy density $\chi_{e}=\delta e/U$ with $\delta e=(1/V)\sum_{\textbf{p}}(\textbf{p}^{2}/2m-\mu)\delta n(\textbf{p})$.
In the hydrodynamic limit, only these conserved quantities remain of order unity, while all the other components pick up one or several factors $\omega_{0}\tau$. To evaluate the power of a given component $\nu_{n}^{l}$ in $\omega_{0}\tau$, one should count the number of transport equations needed to reach a conserved quantity using the couplings appearing in Eq. (20), that is, $\nu_{n}^{l}\to\nu_{n}^{l\pm 1},\nu_{n^{\prime}}^{l}$ (with $n^{\prime}$ having the same parity as $n$). The components $\nu_{n}^{2}$ with $n$ even and $\nu_{n}^{1}$ with $n$ odd, which are directly coupled to the conserved quantities $\nu_{0}^{1}$ and $\nu_{1}^{0}$ respectively, are of order $O(\omega_{0}\tau)^{1}$, and the other components are subleading, as depicted by Fig. 1. In particular, the large $l$ components decay exponentially as $O(\omega_{0}\tau)^{l-1}$ or $O(\omega_{0}\tau)^{l}$, depending on the parity of $n$.
[…]
Figure 1: Schematic of the order in $\omega_{0}\tau$ of the non conserved components $\nu_{n}^{l}$ of the quasiparticle distribution. The red arrows represent the couplings to the conserved quantities $\nu_{0}^{1}$ and $\nu_{1}^{0}$.
The drive on the right-hand-side of Eq. (20) is coupled to $\nu_{0}^{1}$. The perturbation it generates is therefore symmetric in $\epsilon$, that is, $\nu_{n}^{l}=0$ for all $n$ odd. In particular the sound wave does not generate fluctuations of the energy density: $\nu_{1}^{0}=0$. To leading order in $\omega_{0}\tau$ the system Eqs. (22)–(23) describes an ideal hydrodynamic behavior, that is, an undamped resonance at $\omega_{q}=c_{1}q$ with the first sound velocity
$\frac{c_{1}}{v_{\textrm{F}}}=\sqrt{\frac{1+\frac{2k_{\textrm{F}}a}{\pi}}{3}}.$ (25)
To study how the resonance deviates from $c_{1}$, one must compute the set $\vec{\nu}^{2}=(\nu_{n}^{2})_{n}$ of the non-conserved components in the $l=2$ subspace. Keeping the leading and subleading terms in this equation, we obtain
$\left(c+\frac{\textrm{i}}{\pi\omega_{0}\tau}\mathcal{M}^{2}\right)\vec{\nu}^{2}=\frac{2}{3}\nu_{0}^{1}\vec{u}_{0}+O(\omega_{0}\tau)^{2}$ (26)
where we have introduced the matrix $\mathcal{M}^{2}=(\mathcal{M}_{nn^{\prime}}^{l=2})_{nn^{\prime}}$ and the unit vectors $(\vec{u}_{n})_{n^{\prime}}=\delta_{nn^{\prime}}$. We have neglected in Eq. (26) the vector $\vec{\nu}^{3}$ and the components of $\vec{\nu}^{1}$ orthogonal to the zero eigenvector $\nu_{0}^{1}\vec{u}_{0}$: both are of order $O(\omega_{0}\tau)^{2}$ (see Fig. 1), and hence negligible compared to the subleading term $c\vec{\nu}^{2}$ in Eq. (26). As is noted in Ref. Brooker and Sykes (1970), the expansion in powers of $\omega_{0}\tau$ becomes very tedious as soon as higher Legendre components have to be taken into account. Fortunately this is not the case in our calculation to order $O(\omega_{0}\tau)^{2}$: the transport equation can be truncated to $l=2$, and only two collision times are needed to express $\nu_{0}^{2}$ from Eq. (26):
$\nu_{0}^{2}=-\frac{2\textrm{i}\omega_{0}\pi}{3}\nu_{0}^{1}\left[\tau_{\eta}+\textrm{i}\omega_{0}\pi ct_{\eta}^{2}\right]+O(\omega_{0}\tau)^{3}$ (27)
where we introduce the viscous collision time $\tau_{\eta}/\tau=\vec{u}_{0}\frac{1}{\mathcal{M}^{2}}\vec{u}_{0}$ and the second-order viscous time $t_{\eta}/\tau=\sqrt{\vec{u}_{0}\frac{1}{(\mathcal{M}^{2})^{2}}\vec{u}_{0}}$. These two parameters characterize the exact equations of motion of the density and parallel velocity within second-order hydrodynamics; a relaxation time approximation would not distinguish them since it amounts to replacing $\mathcal{M}^{2}$ by a number. In an exact calculation, they can be expressed as infinite series Brooker and Sykes (1968); Højgård Jensen et al. (1968) or as continued fractions Hofmann and Das Sarma (2022), which converge rather slowly when truncated at $n=n^{\prime}=n_{\rm max}$: $\tau_{\eta}-\tau_{\eta}^{(n_{\rm max})}=O(1/n_{\rm max}^{2})$. Numerically, we obtain
$\tau_{\eta}\simeq 1.079\tau_{\eta}^{(0)}\simeq 0.102\tau\text{\ \ and\ \ }t_{\eta}\simeq 1.098\tau_{\eta}^{(0)}\simeq 0.104\tau$ (28)
where we have introduced the viscous collision time truncated at $n=0$: $\tau_{\eta}^{(0)}/\tau=1/\mathcal{M}^{l=2}_{00}=15/16\pi^{2}$.
Figure 2: The sound attenuation and dispersion coefficients $\gamma$ and $\Delta c$ obtained by fitting a Lorentzian function $c\mapsto Z/\left[(c-\bar{c}_{1}-(\omega_{0}\tau)^{2}B(\omega_{0}\tau))^{2}+(\omega_{0}\tau)^{2}A(\omega_{0}\tau)^{2}\right]$ (green and orange lines in inset, respectively for $\omega_{0}\tau=0.04$ and $\omega_{0}\tau=2$) to the density response $\nu_{0}^{0}$ obtained by numerically solving Eq. (20) at non vanishing $\omega_{0}\tau$. The horizontal dashed and dotted line show respectively the exact hydrodynamic limits $B\to-\textrm{Im}(z_{1})/(\omega_{0}\tau)$ and $A\to\textrm{Re}(z_{1}-\bar{c}_{1})/(\omega_{0}\tau)^{2}$ from Eq. (30), and the relaxation time approximation ($\tau_{\eta},t_{\eta}\to\tau_{\eta}^{(0)}$). This figure is drawn in the limit $k_{\textrm{F}}a\to 0$ ($c_{1}/v_{\textrm{F}}=1/\sqrt{3}$).
Plugging Eq. (27) in Eqs. (22)–(23), we obtain a quasi-Lorentzian shape for the density response:
$\chi_{\rho}=\frac{mk_{\textrm{F}}}{6\pi^{2}}\frac{1}{c^{2}-\bar{c}_{1}^{2}+\frac{4\textrm{i}\omega_{0}\pi c}{15}\left(\tau_{\eta}+\textrm{i}\omega_{0}\pi ct_{\eta}^{2}\right)}.$ (29)
The pole $z_{1}$ of this Lorentzian is located at:
$z_{1}=\bar{c}_{1}-\frac{2\textrm{i}\pi}{15}\omega_{0}\tau_{\eta}+\frac{2\pi^{2}\omega_{0}^{2}}{225\bar{c}_{1}}(15\bar{c}_{1}^{2}t_{\eta}^{2}-\tau_{\eta}^{2})$ (30)
where $\bar{c}_{1}=c_{1}/v_{\textrm{F}}$. Going back to the physical units, we write the resonance frequency $\omega_{q}=\text{Re}(z_{1})\omega_{0}$, the main result of this work, as
$\omega_{q}\!=\!c_{1}q\left[1+\theta(\bar{c}_{1})\epsilon_{\textrm{F}}^{2}\tau^{2}\left(\frac{q}{mc_{1}}\right)^{2}\!\!\!+O(\omega_{0}\tau)^{3}\right]$ (31)
with $\theta(c)\simeq 0.057c^{2}-0.0037$. The first deviation from the linear spectrum $c_{1}q$ is thus proportional to $q^{3}$. The dispersion is positive at weak coupling since $\theta(\bar{c}_{1})\simeq 0.0153$ for $\bar{c}_{1}=1/\sqrt{3}$. We note that $\theta(c)$ is negative for $c<c_{\rm inv}\simeq 0.25$, such that an inversion of the sign of the dispersion may occur in settings where the ratio $c_{1}/v_{\textrm{F}}$ is lower than $c_{\rm inv}$.
The damping rate $\Gamma_{q}=-\text{Im}(z_{1})\omega_{0}$ of the sound wave is determined only by the shear viscosity $\eta$ in this low-temperature regime:
$\Gamma_{q}=\frac{2}{3m\rho}\eta q^{2}\text{ with }\eta=\frac{2\pi}{5}\rho\epsilon_{\textrm{F}}\tau_{\eta}\simeq 0.129\rho\epsilon_{\textrm{F}}\tau.$ (32)
This exact value of $\eta$ recovers the result of Refs. Brooker and Sykes (1968); Højgård Jensen et al. (1968); Baym and Pethick (1991) for the case of an isotropic scattering amplitude (see Eq. (11) in Højgård Jensen et al. (1968) with $\alpha=2/5$). This is a factor $\tau_{\eta}/{\tau_{\eta}^{(0)}}\simeq 1.08$ above the value in the relaxation time approximation of Refs. Massignan et al. (2005); Bruun and Smith (2007); Braby et al. (2010) (called the “variational approximation” therein). We note that the underestimation consecutive to the approximation is larger than stated before Bruun and Smith (2007); Braby et al. (2010). Conversely, in the Abrikosov-Khalatnikov approximation Khalatnikov and Abrikosov (110); Nozières and Pines (1966); Watabe et al. (2010), the viscosity is overestimated by a factor ${\tau_{\eta}^{\rm(AK)}}/{\tau_{\eta}}=4\tau_{\eta}^{(0)}/3{\tau_{\eta}}\simeq 1.24$.
Experimentally, the attenuation and dispersion coefficients appearing in Eqs. (31)–(32) can be measured through the shape of the response functions. Fig. 2 shows that they can be accurately recovered from Lorentzian fits of the density response. The shift $\omega_{q}-c_{1}q$ remains positive throughout the crossover from the hydrodynamic to the collisionless regime, as visible on Fig. 3.
Figure 3: The transition from hydrodynamic (red curves) to collisionless regime (blue curves) in the velocity response function $\nu_{n=0}^{l=1}$ of a weakly interacting Fermi gas with $k_{\textrm{F}}a=-0.5$. The broadening and the positive shift of the resonance frequency from the first sound velocity (vertical dashed line, Eq. (25)) is clearly visible at small $\omega_{0}\tau$. The curve are obtained by numerically solving Eq. (20) truncated to $n_{\rm max}=100$ and $l_{\rm max}$ ranging from 20 in the hydrodynamic regime to 8500 in the collisionless regime.
In passing, the equation on the energy density Eq. (24) allows us to compute the thermal diffusivity of the gas. By inverting the matrix $\mathcal{M}^{1}$ in the subspace of the $\nu_{n}^{1}$ with $n$ odd, where it has no conserved quantities, we derive
$\nu_{1}^{1}=-\textrm{i}\pi\omega_{0}\tau_{\kappa}\nu_{1}^{0}\implies\left(\omega+\textrm{i}{D_{\kappa}}q^{2}\right)\nu_{1}^{0}=0\\ \text{ with }D_{\kappa}=\frac{\pi}{3}{v_{\textrm{F}}^{2}}\tau_{\kappa}\simeq 0.0623{v_{\textrm{F}}^{2}\tau}.$ (33)
We have introduced the thermal diffusion time $\tau_{\kappa}/\tau=\vec{u}_{1}\frac{1}{\mathcal{M}^{1}}\vec{u}_{1}$. The relaxation time approximation of this time $\tau_{\kappa}^{(0)}/\tau=15/32\pi^{2}$ is much worse than for the viscous time $\tau_{\eta}$, with the error reaching $25\%$.
Finally, we note that the nonvanishing isotropic components of the collision kernel $\mathcal{M}^{0,\perp}=(\mathcal{M}^{0})_{n>1,n^{\prime}>1}$, which are usually associated with the bulk viscosity $\zeta$, do not affect the propagation of sound, even within second-order hydrodynamics. The decoupling is a consequence of the parabolic dispersion relation (2) of the Fermi liquid quasiparticles, since this does not couple the total velocity $\nu_{0}^{1}$ directly to the nonconserved isotropic components of the quasiparticle distribution $\nu_{n>1}^{0}$. This means the bulk viscosity appears in the hydrodynamic equations only at higher temperatures and stronger interactions. Nevertheless, the comparison of the exact to approximate matrix elements $(1/\mathcal{M}^{0,\perp})_{22}\simeq 1.41/\mathcal{M}^{0}_{22}$ suggests that the error made by the relaxation time approximations on the value of the bulk viscosity is large.
Acknowledgements.
H.K. acknowledges support from the EUR grant NanoX n°ANR-17-EURE-0009 in the framework of the “Programme des Investissements d’Avenir”. N.N acknowledges support by the AFOSR (Grant No. FA9550-23-1-0605), NSF (Grant No. PHY-1945324), DARPA (Grant No. HR00112320038), and the David and Lucile Packard Foundation. T.R. and H.K. thank Yale University for its hospitality.